### v1

@@ Abstract
The use of stochastic RCs holds significant potential because the reservoir can be designed such that the number of computational nodes and, therefore, the possible outcomes scale exponentially with the hardware size. For example, in a qubit-based quantum computer, one can perform a unitary rotation of a set of $n_{q}$ qubits based on the input at a given time step, then measure the qubits in the computational basis. By running this process many times, we can obtain an estimate of the probability of measuring each of the $2^{n_{q}}$ distinct bit strings and use them as readouts for training to get an output. The requirement that the stochastic RC must be run many times to get an estimate of the probabilities is a potential setback, representing a tradeoff between more computational nodes and machine runtime. However, this approach becomes valuable in scenarios where a deterministic computer lacks sufficient computational nodes to perform the necessary calculations. In such cases, a stochastic machine can achieve an equivalent number of computational nodes using a much smaller device. We examine two different designs for a stochastic RC more closely in the Results section, where we define the exponential scaling more concretely.
@@
The reason this construction is useful is because it is easy to achieve a large number of outcomes $M$ by stringing together a small number of stochastic systems with only a few outcomes individually and reading the outcomes out as a vector $x_{k}$ at each time step. For example, a binary stochastic process (a $p$-bit [24]) has only two possible outcomes, but $L$ of these together have $M=2^{L}$ distinct outcomes, and in general a vector outcome $x$ whose components each have $m$ possible outcomes will have as an ensemble $M=m^{L}$ distinct outcomes. This makes it easy to generate a RC with a large number of computational nodes using a fairly small device, as we will illustrate in the Results section.
@@
This can be written more succinctly in matrix form as $P_{k+1}=p(u_{k})P_{k}$. This means that the probabilities themselves form a Markov chain [25], in particular a controlled discrete-time Markov chain [26] in which the input sequence $(u_{k})_{k\in\mathbb{Z}}=\{u_{k}\left|k\in\mathbb{Z}\right.\}$ acts as the fixed control sequence for the Markov chain. The controlled transition probability matrix $p(u)$ whose elements are defined by (4) is determined by the dynamical reservoir. This means that, rather than using the finite number of outcomes in the set $\mathcal{X}$ as readouts, we can instead use the probability vectors $P_{k}$ as our readouts and train those to fit a target series. Thus the controlled Markov chain forms the basis of stochastic reservoir computing, and we can then write the RC equations as
@@
where $\hat{y}_{k}$ is our estimate for some target value $y_{k}$ which is fit by training the $M$-dimensional weight vector $W$.
@@ Results
In the next couple of sections we will demonstrate the convergence and performance of stochastic ESNs using a couple of examples. In both examples, we use a stochastic activation function $\Gamma(\zeta)$ that outputs either 00 or $1$ with controlled probabilities $\varrho_{0}(\zeta)$ and $\varrho_{1}(\zeta)=1-\varrho_{0}(\zeta)$ that have the required properties which ensures that the class $\mathcal{G}_{w}^{\varrho}$ of stochastic ESNs using $\varrho(\zeta)$ is a universal approximating class. The designs of the two examples are also meant to emphasize the practicality of implementing a stochastic RC using readily available hardware. In particular, they can be scaled up in a relatively easy way to produce an exponentially scaling number of computational nodes, with the drawback of potentially needing a larger number of measurements to estimate the probabilities to the required level of precision.
To test the performance of these stochastic ESNs, we will use two different tasks, the first of which is the Sine-Square wave identification task [27]. This task requires the ESN to identify whether a given section of the input is a sine wave or a square wave, yielding a output of $\hat{y}_{k}<0.5$ for a sine wave and $\hat{y}_{k}>0.5$ for a square wave. The other task we will use in this work is the Lorenz $X$ task. The Lorenz $X$ task requires the ESN to approximate a chaotic dynamical system and predict what the sequence will be 1 time step into the future. In other words, using the input sequence $(u_{k})_{k\in\mathbb{Z}}=(y_{k-1})_{k\in\mathbb{Z}}$, we want the ESN to successfully predict $(y_{k})_{k\in\mathbb{Z}}$ in this task. Further details about these tasks and the simulation parameters are given in Supplementary Information A.3.
@@
The qubit reservoir network is an example of a stochastic echo state network that utilizes simple, single-qubit gate operations applied to an array of qubits to generate a nonlinear stochastic activation function. This is a new hybrid quantum-classical approach that capitalizes on the ease of performing linear operators on classical computers while also leveraging a quantum device for rapid nonlinear transformations, similar to a variational quantum algorithm. The routine for each time step of the ESN is illustrated in Fig. 1, where the linear transformation is applied classically, and the $z$ values are passed onto each qubit using a transformation operator of the form $U(\theta)=e^{-i\sigma_{x}\theta}$. First, at each time step $k$ we use the most recent array of qubit measurements $x_{k}$ and the current input $u_{k}$ to calculate $z_{k}=Ax_{k}+Bu_{k}$ on a classical computer. Then, we apply the transformation operator $U(z_{k,i}+\frac{\pi}{2}x_{k,i}+\delta)$ to each qubit, where $\delta$ is some constant shift parameter. The extra factor of $\frac{\pi}{2}x_{k,i}$ resets the qubits to $|0\rangle$ from the states $|x_{k,i}\rangle$ they collapsed to after the previous measurement. Finally, each qubit is then measured in the computational basis, which is recorded as $x_{k+1}$ and used immediately in the next time step. We then repeat this cycle many times and record the number of times each distinct measurement outcome $x^{(a)}$ appears at each time step $k$ to obtain estimates of their probabilities $P_{k,a}$ of occurring. These probabilities are then used as the readout vector for training and obtaining estimates $\hat{y}_{k}=W^{\top}P_{k}$ of the target output $y_{k}$.
@@
This class of stochastic ESNs can be shown to be universal when $\delta=\frac{\pi}{4}$ and when the matrix $A$ and vector $B$ are chosen so that all elements of $z_{k}=Ax_{k}+Bu_{k}$ satisfy $|z_{k,i}|<\frac{\pi}{4}$ at every time step $k$. We show this in Supplementary Information A.1.
Refer to captionRefer to caption
@@
The left plot of Fig. 2 shows a comparison of NMSE values for the Lorenz $X$ task averaged over 100 different choices of $A$ matrices and $B$ vectors for the ESN, while the right plot shows a comparison of the percentage of errors recorded for the Sine-Square wave identification task averaged over 1000 choices. We see in the figure that the stochastic ESN using the exact probabilities performs much better than the deterministic ESN using roughly the same hardware. With finite statistics, the stochastic ESN still performs well for a small number of detectors, but beyond 4 detectors the values of the error measure seems to level off, and the deterministic ESN begins to outperform it on average by 4 detectors in the Lorenz $X$ task and at 7 detectors in the Sine-Square wave task.
An explanation for this can be found by analysing how the linear regression behaves under the noise introduced by the stochastic reservoir, which is detailed in Supplementary Information A.2. The takeaway is that if $N_{\mathrm{runs}}$ is not large enough to suppress the shot noise represented by $P_{k}^{\mathrm{diag}}$ relative to all of the eigenvalues of $K$, then the performance of the stochastic reservoir will be reduced compared to what it could be with the exact probabilities.
@@
This class of stochastic ESNs can be shown to be universal with $\delta=1$ and when the matrix $A$ and vector $B$ are chosen so that all elements of $z_{k}=Ax_{k}+Bu_{k}$ satisfy $|z_{k,i}|<1$ at every time step $k$. We show this in Supplementary Information A.1.
Refer to captionRefer to caption
@@ Discussion
Looking more closely at Figs. 2 and 4, it appears that the error measures in both cases for the stochastic ESNs using the exact probabilities also level off after a certain point, much like the 10000 run ESNs as discussed in Supplementary Information A.2. In fact, if we keep increasing the number of detectors of the deterministic ESNs beyond the values shown in the figures, we would see the same leveling off of performance as well. Considering that the smallest eigenvalue of $K$ continues getting smaller and smaller as the number of detectors rises, it seems as though machine precision errors may substantially limit the performance of larger RCs. One way to potentially avoid this would be to use an alternative to optimizing via matrix inversion, such as gradient descent. However, gradient descent methods have their own issues resolving small values [33, 34, 35] which would likely be present in this context. Another option would be to find reservoirs that do not lead to small eigenvalues in $K$. These RCs would be able to achieve a manageable signal-to-noise ratio for all of its eigentasks (see Ref. [36]), but this would require a much stronger theoretical handle on how reservoir dynamics lead to strong fits, which is at present a topic in need of extensive research.
@@ Methods
Our proof of universality for stochastic RCs will follow the strategy used in Refs. [13, 16], proving that the class of stochastic RCs satisfies the criterion of the Stone-Weierstrass Theorem [37, Theorem 7.3.1]. In the next few sections, we will explicitly prove the universality of certain classes of stochastic ESNs, which implicitly proves the universality of the class containing all stochastic RCs with the contraction property because ESNs are a subclass of RCs.
@@
Since convergence criterion of the conditional probability matrix is a well-known result, the contents of Thm. 3 are almost surely in previous literature, and we do not provide a proof in this section. For completeness, we do provide a proof in Supplementary Information A.4.
@@
With polynomial algebra proven and sufficient criteria for separability given for classes of stochastic ESNs, we are now finally ready to apply the Stone-Weierstrass Theorem [37, Theorem 7.3.1] and prove the universality of these classes.
@@
Appendix A Supplementary Information
A.1 Universality of Proposed Designs
@@
A.2 Noise Analysis
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ A.2 Noise Analysis
A.3 Simulation Details
A.3.1 Task Details
@@ A.3.2 Simulation Parameters
For each run, we used 3000 training data steps performed after 100 steps of initial startup time for the Lorenz $X$ task and 96 steps for the Sine-Square wave task. The values for the error measures shown here are for a set of time steps immediately after the training data set, totaling 500 steps for the Lorenz $X$ task and 504 for the Sine-Square wave task. These same parameters apply to the data in Figs. 5 and 6 from the previous section as well. However, for the Sine-Square wave plots in these figures, each data point is an average over 10 repetitions of the error percentage calculation, using the same number of runs each time, to reduce the spread of the values in this plot.
A.4 Proof of Theorem 3
@@
A.5 Note on a Recent Result
Stochastic RCs as we define them have been considered in a recent paper by Polloreno [41]. In this work, it is shown that the information processing capacity of a single run of a stochastic RC only scales polynomially with the scale of the hardware. However, in our work we run the RC many times and get estimates for the probabilities $P_{k,a}$ by counting the number of occurrences of each outcome. The derivation of Theorem 2 in Ref. [41] is predicated on the fact that the second moment of the outcome probabilities $P_{k,a}$ is linear in these probabilities, but this is no longer true when $P_{k,a}$ is estimated with an average over multiple shots. In Supplementary Information A.2, we show that past a certain number of aggregated runs, the estimated probabilities provide a very good estimate of the results using the exact probabilities. Of course, it could still be argued that the number of runs needed to achieve this level of accuracy may be exponentially scaling with the scale of the hardware as well. In any case, we do not claim that stochastic RCs have exponentially scaling computational complexity. We merely state that the number of computational nodes of the reservoir is exponentially scaling, and these nodes may or may not be used effectively for any given computation.

### v2

@@ Abstract
The use of stochastic RCs holds significant potential because the reservoir can be designed such that the number of computational nodes and, therefore, the possible outcomes scale exponentially with the hardware size. For example, in a qubit-based quantum computer, one can perform a unitary rotation of a set of $n_{q}$ qubits based on the input at a given time step, then measure the qubits in the computational basis. By running this process many times, we can obtain an estimate of the probability of measuring each of the $2^{n_{q}}$ distinct bit strings and use them as readouts for training to get an output. Another potentially promising implementation of stochastic RCs is on recently proposed classical (non-quantum) probabilistic computers that are based on probabilistic bits (p-bits) and other classical stochastic devices [24, 25, 26], including classical devices that exploit quantum effects [25]. The requirement that the stochastic RC must be run many times to get an estimate of the probabilities is a potential setback, representing a tradeoff between more computational nodes and machine runtime. However, this approach becomes valuable in scenarios where a deterministic computer lacks sufficient computational nodes to perform the necessary calculations. In such cases, a stochastic machine can achieve an equivalent number of computational nodes using a much smaller device. We examine two different designs for a stochastic RC more closely in the Results section, where we define the exponential scaling more concretely.
@@
The reason this construction is useful is because it is easy to achieve a large number of outcomes $M$ by stringing together a small number of stochastic systems with only a few outcomes individually and reading the outcomes out as a vector $x_{k}$ at each time step. For example, a binary stochastic process (a p-bit) has only two possible outcomes, but $L$ of these together have $M=2^{L}$ distinct outcomes, and in general a vector outcome $x$ whose components each have $m$ possible outcomes will have as an ensemble $M=m^{L}$ distinct outcomes. This makes it easy to generate a RC with a large number of computational nodes using a fairly small device, as we will illustrate in the Results section.
@@
This can be written more succinctly in matrix form as $P_{k+1}=p(u_{k})P_{k}$. Therefore, any stochastic RC as defined here is a Markov chain [27], in particular a controlled discrete-time Markov chain [28] in which the input sequence $(u_{k})_{k\in\mathbb{Z}}=\{u_{k}\left|k\in\mathbb{Z}\right.\}$ acts as the fixed control sequence for the controlled Markov chain. The controlled transition probability matrix $p(u)$ whose elements are defined by (4) is determined by the dynamical reservoir. This means that, rather than using the finite number of outcomes in the set $\mathcal{X}$ as readouts, we can instead use the probability vectors $P_{k}$ as our readouts and train those to fit a target series. Thus the controlled Markov chain forms the basis of stochastic reservoir computing, and we can then write the RC equations as
@@
where $\hat{y}_{k}$ is our estimate for some target value $y_{k}$ which is fit by training the $M$-dimensional weight vector $W$. This causes $P_{k}$ and $\hat{y}_{k}$ to have dependence on all past inputs $\{u_{k-1},u_{k-2},\dots\}$ while remaining independent of the current input $u_{k}$. The reasoning for this is that physical systems do not have an instantaneous response to their environment, so in modeling the functional dependence of such a system with a RC we should also avoid using the current value of the input.
@@ Results
In the next couple of sections we will demonstrate the convergence and performance of stochastic ESNs using a couple of examples. In both examples, we use a stochastic activation function $\Gamma(\zeta)$ that outputs either 00 or $1$ with controlled probabilities $\varrho_{0}(\zeta)$ and $\varrho_{1}(\zeta)=1-\varrho_{0}(\zeta)$ that have the required properties which ensures that the class $\mathcal{G}_{w}^{\varrho}$ of stochastic ESNs using $\varrho(\zeta)$ is a universal approximating class. A precise statement of these properties is given at the top of the Methods section. The designs of the two examples are also meant to emphasize the practicality of implementing a stochastic RC using readily available hardware. In particular, they can be scaled up in a relatively easy way to produce an exponentially scaling number of computational nodes, with the drawback of potentially needing a larger number of measurements to estimate the probabilities to the required level of precision.
To test the performance of these stochastic ESNs, we will use two different tasks, the first of which is the Sine-Square wave identification task [29]. This task requires the ESN to identify whether a given section of the input is a sine wave or a square wave, yielding a output of $\hat{y}_{k}<0.5$ for a sine wave and $\hat{y}_{k}>0.5$ for a square wave. The other task we will use in this work is the Lorenz $X$ task. The Lorenz $X$ task requires the ESN to approximate a chaotic dynamical system and predict what the sequence will be 1 time step into the future. In other words, using the input sequence $(u_{k})_{k\in\mathbb{Z}}=(y_{k-1})_{k\in\mathbb{Z}}$, we want the ESN to successfully predict $(y_{k})_{k\in\mathbb{Z}}$ in this task. In both tasks, we run the ESN for roughly 100 steps so that the system will forget its initial state, then train the ESN using 3000 steps, then test the performance on roughly 100 time steps afterward. Further details about these tasks and the simulation parameters are given in Supplementary Information S.3. The data and code used to obtain our results can be found in a figshare repository [30].
@@
The qubit reservoir network is an example of a stochastic echo state network that utilizes simple, single-qubit gate operations applied to an array of qubits to generate a nonlinear stochastic activation function. This is a new hybrid quantum-classical approach that capitalizes on the ease of performing linear operators on classical computers while also leveraging a quantum device for rapid nonlinear transformations, similar to a variational quantum algorithm. The routine for each time step of the ESN is illustrated in Fig. 1, where the linear transformation is applied classically, and the $z$ values are passed onto each qubit using a transformation operator of the form $U(\theta)=e^{-i\sigma_{x}\theta}$. First, at each time step $k$ we use the most recent array of qubit measurements $x_{k}$ and the current input $u_{k}$ to calculate $z_{k}=Ax_{k}+Bu_{k}$ on a classical computer. Then, we apply the transformation operator $U(z_{k,i}+\frac{\pi}{2}x_{k,i}+\delta)$ to each qubit, where $\delta$ is some constant shift parameter. The extra factor of $\frac{\pi}{2}x_{k,i}$ resets the qubits to $|0\rangle$ from the states $|x_{k,i}\rangle$ they collapsed to after the previous measurement. Finally, each qubit is then measured in the computational basis, which is recorded as $x_{k+1}$ and used immediately in the next time step. We then repeat this cycle many times and record the number of times each distinct measurement outcome $x^{(a)}$ appears at each time step $k$ to obtain estimates of their probabilities $P_{k,a}$ of occurring. These probabilities are then used as the readout vector for training and obtaining estimates $\hat{y}_{k}=W^{\top}P_{k}$ of the target output $y_{k}$. The initial state of the qubits has a negligible effect on the results because the ESN necessarily has the uniform convergence and fading memory properties, which are defined in the Methods section. These properties make it so that the initial condition will be forgotten as the reservoir is driven to some specific evolving configuration by the inputs. By letting the system run for a number of time steps before training, the dependence on the initial condition is effectively removed.
@@
This class of stochastic ESNs can be shown to be universal when $\delta=\frac{\pi}{4}$ and when the matrix $A$ and vector $B$ are chosen so that all elements of $z_{k}=Ax_{k}+Bu_{k}$ satisfy $|z_{k,i}|<\frac{\pi}{4}$ at every time step $k$. We show this in Supplementary Information S.1. The value of $\delta=\frac{\pi}{4}$ is chosen because it allows for the largest possible upper bound for $|z_{k,i}|$ that still leaves $\varrho_{1}(\zeta)$ monotonic and therefore generates a universal approximating class. Other choices can still achieve universality, but will require tighter bounds on $|z_{k,i}|$ and thus better numerical precision. One could use an ESN that goes outside of the bound on $|z_{k,i}|$, and it may even perform well, but there is no guarantee that a useful ESN exists outside this window.
@@
The left plot of Fig. 2 shows a comparison of NMSE values for the Lorenz $X$ task averaged over 100 different choices of $A$ matrices and $B$ vectors for the ESN, while the right plot shows a comparison of the percentage of errors recorded for the Sine-Square wave identification task averaged over 1000 choices. In this context, “# of detectors” refers to the number of physical qubits that make up the reservoir. We see in the figure that the stochastic ESN using the exact probabilities performs much better than the deterministic ESN using roughly the same hardware. With finite statistics, the stochastic ESN still performs well for a small number of detectors, but beyond 4 detectors the values of the error measure seems to level off, and the deterministic ESN begins to outperform it on average by 4 detectors in the Lorenz $X$ task and at 7 detectors in the Sine-Square wave task.
An explanation for this can be found by analysing how the linear regression behaves under the noise introduced by the stochastic reservoir, which is detailed in Supplementary Information S.2. The takeaway is that if $N_{\mathrm{runs}}$ is not large enough to suppress the shot noise and resolve all of the basis vectors formed from the $P_{k}$’s, then the performance of the stochastic reservoir will be reduced compared to what it could be with the exact probabilities.
@@
This class of stochastic ESNs can be shown to be universal with $\delta=1$ and when the matrix $A$ and vector $B$ are chosen so that all elements of $z_{k}=Ax_{k}+Bu_{k}$ satisfy $|z_{k,i}|<1$ at every time step $k$. We show this in Supplementary Information S.1.
@@ Discussion
Looking more closely at Figs. 2 and 4, it appears that the error measures in both cases for the stochastic ESNs using the exact probabilities also level off after a certain point, much like the 10000 run ESNs as discussed in Supplementary Information S.2. In fact, if we keep increasing the number of detectors of the deterministic ESNs beyond the values shown in the figures, we would see the same leveling off of performance as well. Considering that adding more nodes will introduce basis vectors that always seem to get more difficult to resolve, it seems as though machine precision errors may substantially limit the performance of larger RCs. For further discussion on this point, see the end of section S.2 in the Supplementary Information. One way to potentially avoid this would be to use an alternative to optimizing via matrix inversion, such as gradient descent. However, gradient descent methods have their own issues resolving small values [36, 37, 38] which would likely be present in this context. Another option would be to find reservoirs with basis vectors that are easy to resolve. These RCs would be able to achieve a manageable signal-to-noise ratio for all of its eigentasks (see Ref. [39]), but this would require a much stronger theoretical handle on how reservoir dynamics lead to strong fits, which is at present a topic in need of extensive research.
@@ Methods
Our proof of universality for stochastic RCs will follow the strategy used in Refs. [13, 16], proving that the class of stochastic RCs satisfies the criterion of the Stone-Weierstrass Theorem [40, Theorem 7.3.1]. In the next few sections, we will explicitly prove the universality of certain classes of stochastic ESNs, which implicitly proves the universality of the class containing all stochastic RCs with the contraction property because ESNs are a subclass of RCs.
@@
The conditions that guarantee the universality of the associated approximating class are given here as reference. The reasons for asserting these conditions will be elaborated on in the proofs below. In terms of the probability distribution vector $\varrho(\zeta)$ defined on some interval $\zeta\in[-R_{\zeta},R_{\zeta}]$ with $R_{\zeta}>0$, these conditions can be written as follows:
  1. 1.
    $\varrho(\zeta)$ is continuous in $\zeta$ for all $\zeta\in[-R_{\zeta},R_{\zeta}]$.
  2. 2.
    $\varrho(\zeta)$ satisfies $\varrho^{\top}(\zeta_{1})\varrho(\zeta_{2})>0$ for all $\zeta_{1},\zeta_{2}\in[-R_{\zeta},R_{\zeta}]$.
  3. 3.
    For the invertible subspace $\mathcal{V}_{R_{u}}^{p}$ defined in the Methods section, there exists a vector $V\in\mathcal{V}_{R_{u}}^{p}$ for which $V^{\top}\varrho(\zeta)$ is strictly monotonic on the interval $[-R_{\mathrm{mono}},R_{\mathrm{mono}}]$ for some $R_{\mathrm{mono}}>0$.
The first two criteria ensure that the reservoir computer has the uniform convergence and fading memory properties. More specifically, criterion 2 guarantees that the action of the reservoir is contracting in the probabilities over time, which in turn guarantees uniform convergence. This combined with the continuity of $\varrho(\zeta)$ also leads to the fading memory property. Criterion 3 is required in order to separate points in the space of input histories, which is required for a universal approximating class. The monotonicity of the vector component $V^{\top}\varrho(\zeta)$ ensures that any difference between inputs can be resolved, no matter how small. The requirement that the vector $V$ belongs to the invertible subspace is necessary to make sure that some vector $W$ exists that can pick out the monotonic component of $\varrho(\zeta)$ at any point in the past, which is necessary to prove separability of two input sequences that are identical after some time step in the past.
In the special case where $\varrho(\zeta)=\left(\begin{matrix}\varrho_{1}(\zeta)&(1-\varrho_{1}(\zeta))\end{matrix}\right)^{\top}$, which applies to both numerical examples described earlier, the above conditions can be rewritten in terms of the scalar function $\varrho_{1}(\zeta)$, and will reduce to the following:
  1. 1.
    $\varrho_{1}(\zeta)$ is continuous in $\zeta$ for all $\zeta\in[-R_{\zeta},R_{\zeta}]$.
  2. 2.
    either $\varrho_{1}(\zeta)\neq 0$ or $\varrho_{1}(\zeta)\neq 1$ for all $\zeta\in[-R_{\zeta},R_{\zeta}]$.
  3. 3.
    $\varrho_{1}(\zeta)$ is strictly monotonic on the interval $[-R_{\mathrm{mono}},R_{\mathrm{mono}}]$ for some $R_{\mathrm{mono}}>0$.
The second statement says that Criterion 2 is automatically satisfied when the chances of measuring either outcome are never certain. It also turns out that the invertible subspace requirement in Criterion 3 becomes trivially satisfied when there are only two measurement outcomes per detector, so we can place the monotonicity requirement directly on $\varrho_{1}(\zeta)$ instead. For details on why these are true, see the Corollaries below.
@@
Since convergence criterion of the conditional probability matrix is a well-known result, the contents of Thm. 3 are almost surely in previous literature, and we do not provide a proof in this section. For completeness, we do provide a proof in Supplementary Information S.4.
@@
With polynomial algebra proven and sufficient criteria for separability given for classes of stochastic ESNs, we are now finally ready to apply the Stone-Weierstrass Theorem [40, Theorem 7.3.1] and prove the universality of these classes.
@@
Data and Code Availability
The data generated for the figures in this article as well as the Jupyter notebooks written in Python used to make them are available in a figshare repository https://doi.org/10.6084/m9.figshare.27957141.
Appendix S Supplementary Information
S.1 Universality of Proposed Designs
@@
S.2 Noise Analysis
@@
Figure 7: The left plot shows average NMSE values for the Lorenz $X$ task on the stochastic optical network, while the right plot shows the average error percentage for the Sine-Square wave identification task on the qubit reservoir network, both as a function of the number of runs of the stochastic ESN. The different lines shown correspond to the averages for ESNs with different numbers of detectors, specifically for 2, 3, 4, and 6 detectors. The point labeled $\infty$ on the x-axes of these plots correspond to the stochastic ESN results in the theoretical limit of infinite statistics, using the exact probabilities. Each data point represents the average over 100 random choices of $A$ and $B$, and the error bars give the standard deviations of each data point over the samples of ESNs with different choices of $A$ and $B$.
Figure 8: The left plot shows average NMSE values for the Lorenz $X$ task on the stochastic optical network, while the right plot shows the average error percentage for the Sine-Square wave identification task on the qubit reservoir network, both as a function of the number of computational nodes of the ESNs. The blue points show the NMSE values for the deterministic ESNs, while the orange points show the results for the stochastic ESN in the theoretical limit of infinite statistics, using the exact probabilities. Note that in this plot we are comparing the deterministic and stochastic reservoirs based on the number of computational nodes, rather than by the number of hardware detectors as done in previous figures. Each data point represents the average over 1000 random choices of $A$ and $B$, and the error bars give the standard deviations of each data point over the samples of ESNs with different choices of $A$ and $B$.
In Fig. 7, we show how the average performance of the stochastic ESNs approaches the limit of infinite statistics as the number of runs increases. This differs from Figs. 5 and 6 in that it illustrates the average performance of the stochastic ESNs for $n_{d}\geq 2$ rather than the performance of specific ESNs with $n_{d}=2$. We see that the ESNs generally converge to the infinite run limit by 100000 runs when $n_{d}=2$, but for more detectors there remains a sizable gap in performance between these two points. This necessitates looking into noise mitigation techniques in future work to approach the theoretical limit more quickly for larger ESNs.
In the Discussion section, we observe a level off of the performance in the stochastic reservoirs as the number of detectors increase, and claim that this will also occur in deterministic RCs with more than 8 nodes. We suggest that this is due to the limited machine precision of our Python simulations, which induce errors that substantially limit the performance of larger RCs, even when the RC is otherwise deterministic.
In Fig. 8, we compare the average performance of the deterministic and stochastic RCs in terms of the number of computational nodes rather than the number of detectors as in previous figures. The contrast is in the fact that the previous figures were intended to illustrate the computational capabilities of deterministic and stochastic RCs as a function of their physical size, which is represented by the number of detectors. On the other hand, with Fig. 8 we are now trying to present evidence of a purely computational phenomenon that arises as a result of training the weights associated with the RC, which seems to occur regardless of the physical characteristics of the RC, so a comparison based on the number of computational nodes is most appropriate here. For the deterministic RC, we have $n_{c}=n_{d}+1$ because we have $\hat{y}_{k}=W^{\top}x_{k}+C$, with $n_{d}$ nodes coming from each component of $W$ and an extra node coming from $C$, while for stochastic RCs $n_{c}=2^{n_{d}}$ because $\hat{y}_{k}=W^{\top}P_{k}$. We emphasize again that for the same number of computational nodes $n_{c}$ a stochastic RC can have a smaller physical size (smaller $n_{d}$) than its deterministic counterpart.
Fig. 8 shows the average performance of the deterministic RCs on the two tasks we focus on for up to 50 computational nodes, illustrating the level off claimed earlier. Also included are the data from the stochastic RC calculations in the limit of infinite statistics that were previously shown in Figs. 2 and 4. In the figure, we observe that the deterministic ESN in the Lorenz-X task exhibits diminishing returns after 10 computational nodes, and fails to show any improvement at around 30 nodes. For the number of errors in the Sine-Square task, we see that the best performance occurs at around 15 nodes in the deterministic reservoir, after which both the average and the variance increase with the number of nodes, presumably due to overfitting. This phenomenon is also evident in the stochastic reservoir using the exact probabilities, though for the Sine-Square wave task the best performance seems to occur closer to 32 computational nodes in that case.
Our explanation for why this occurs and how it relates to machine precision is the following. In essence, the components of the readout vectors seem to be too similar to each other, so that the differences between them are orders of magnitude less than the vectors themselves. This means that the optimized training weights must amplify these small differences to get a different output function, which will lead to significant noise contamination if these differences are not resolved. In training, we must invert the matrix $K=\frac{1}{N}\sum_{k}P_{k}P_{k}^{\top}$ for the stochastic ESN, with $P_{k}$ replaced with $x_{k}$ for the deterministic ESNs. As the number of nodes increase in these tasks, we observe that the larger eigenvalues and their corresponding eigenvectors remain largely fixed as the number of nodes increases, while new eigenvectors are added which correspond to smaller and smaller eigenvalues. Eventually, these eigenvalues hit a lower limit based on machine precision where they tend to emerge at roughly the same order of magnitude rather than at lower magnitudes as before, indicating an fundamental change in the fit. We use Tikhonov regularization for the matrix inversion with a regularization parameter of $10^{-10}$, which provides the actual lower limit to these eigenvalues in our simulations and is known to produce an effect similar to noise contamination, but this regularization is also commonplace and necessary to avoid the direct consequences of machine precision errors in RCs with many computational nodes. Due to the wide-reaching implications of such an issue in reservoir computing, we believe that this deserves a more thorough investigation that is outside of the scope of this article.
S.3 Simulation Details
S.3.1 Task Details
@@
For each run, we used 3000 training data steps performed after 100 steps of initial startup time for the Lorenz $X$ task and 96 steps for the Sine-Square wave task. The values for the error measures shown here are for a set of time steps immediately after the training data set, totaling 500 steps for the Lorenz $X$ task and 504 for the Sine-Square wave task. These same parameters apply to the data in all of the figures from the previous section as well. However, for the Sine-Square wave plots in Figs. 5 and 6, each data point is an average over 10 repetitions of the error percentage calculation, using the same number of runs each time, to reduce the spread of the values in this plot.
S.4 Proof of Theorem 3
@@
S.5 Note on a Recent Result
Stochastic RCs as we define them have been considered in a recent paper by Polloreno [44]. In this work, it is shown that the information processing capacity of a single run of a stochastic RC only scales polynomially with the scale of the hardware. However, in our work we run the RC many times and get estimates for the probabilities $P_{k,a}$ by counting the number of occurrences of each outcome. The derivation of Theorem 2 in Ref. [44] is predicated on the fact that the second moment of the outcome probabilities $P_{k,a}$ is linear in these probabilities, but this is no longer true when $P_{k,a}$ is estimated with an average over multiple shots. In Supplementary Information S.2, we show that past a certain number of aggregated runs, the estimated probabilities provide a very good estimate of the results using the exact probabilities. Of course, it could still be argued that the number of runs needed to achieve this level of accuracy may be exponentially scaling with the scale of the hardware as well. In any case, we do not claim that stochastic RCs have exponentially scaling computational complexity. We merely state that the number of computational nodes of the reservoir is exponentially scaling, and these nodes may or may not be used effectively for any given computation.